A new screening test for Lyme disease is developed for use in the general population. The sensitivity and specificity of the new test are 60% and 70%, respectively. Three hundred people are screened at a clinic during the first year the new test is implemented. Assume the true prevalence of Lyme disease among clinic attendees is 10%.

Calculate the following values:

The predictive value of a positive test

The predictive value of a negative test
Answer #1

tinen fo tal number of people =300 Sensitivity = 601. Speci ficity= 70% Sensitivity- TP Specifidty= TN FN+TN -Xl00 TP+FN %3DTP+FN = 30 So 을FN+ Fz = 30 ㅋ FN= 30 - FN=12/ And TP-30- Fu=30-12=/8 TP=1R we Kmoww - TP Since FP+FN = 30, FP+TN = 300-30=270Now (1) Bedi ctive value of positive Test (PPU): 18 Xl00 = 18. 18/8% 99 TPHFP PPV =18.1818%. (2)The Bedictive Value of Negati